# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ Appendix A A Examples of the Weak-strong Uniqueness Classes
This is a direct consequence of [DC06, Theorem 1.1]. We sketch the proof for completeness. For every $t\in[0,T]$, we expand $\|\theta_{1}(t)-\theta_{2}(t)\|_{H}^{2}$ as
$\|\theta_{1}(t)-\theta_{2}(t)\|_{H}^{2}=2\langle\theta_{1}(t),\theta_{2}(t)\rangle+\|\theta_{1}(t)\|_{H}^{2}+\|\theta_{2}(t)\|_{H}^{2}$ (A.10)
and estimate these three terms respectively. For the first term of the right-hand side of (A.10), we need the following auxiliary identity (see [DC06, (2.1)]):
$\quad\int_{0}^{t}\left(2\langle\Lambda^{\alpha}\theta_{1},\Lambda^{\alpha}\theta_{2}\rangle+\langle u_{\theta_{2}}\cdot\nabla\theta_{2},\theta_{1}\rangle+\langle u_{\theta_{1}}\cdot\nabla\theta_{1},\theta_{2}\rangle\right)\,\mathrm{d}s$ (A.11)
$=-\langle\theta_{2}(t),\theta_{1}(t)\rangle+\langle\theta_{2}(0),\theta_{1}(0)\rangle+\int_{0}^{t}\langle\Lambda^{\alpha}(\theta_{2}+\theta_{1}),g\rangle\,\mathrm{d}s.$
Since $\theta_{1}\in\mathcal{C}_{0}^{2}$ is sufficiently regular to be a test function ([Cha06b]), we can apply the $H$-inner product to $\theta_{1}$ to obtain that
$\frac{1}{2}\|\theta_{1}(t)\|_{H}^{2}+\int_{0}^{t}\|\theta_{1}(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\left\|\theta_{1}(0)\right\|_{H}^{2}+\int_{0}^{t}\langle\Lambda^{\alpha}\theta_{1},g\rangle\,\mathrm{d}s.$ (A.12)
According to (A.10), (A.11), (A.12) and (A.9), we obtain
$\|\theta_{1}(t)-\theta_{2}(t)\|_{H}^{2}$ $+\int_{0}^{t}\left\|\theta_{1}(s)-\theta_{2}(s)\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ (A.13)
$\leqslant-\int_{0}^{t}\left\|\theta_{1}(s)-\theta_{2}(s)\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+2\int_{0}^{t}\langle u_{\theta_{2}}\cdot\nabla\theta_{2},\theta_{1}\rangle+\langle u_{\theta_{1}}\cdot\nabla\theta_{1},\theta_{2}\rangle\,\mathrm{d}s$
$=-\int_{0}^{t}\left\|\theta_{1}(s)-\theta_{2}(s)\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+2\int_{0}^{t}\langle(u_{\theta_{2}}-u_{\theta_{1}})\cdot\nabla(\theta_{2}-\theta_{1}),\theta_{1}\rangle\,\mathrm{d}s.$
Similar to (A.8), we have the estimate
$\quad\int_{0}^{t}\langle(u_{\theta_{2}}-u_{\theta_{1}})\cdot\nabla(\theta_{2}-\theta_{1}),\theta_{1}\rangle\,\mathrm{d}s$ (A.14)
$\leqslant C\int_{0}^{t}\|u_{\theta_{2}}-u_{\theta_{1}}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\theta_{2}-\theta_{1}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\,\mathrm{d}s$
$\leqslant C\int_{0}^{t}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\theta_{2}-\theta_{1}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
$\leqslant C\int_{0}^{t}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\theta_{2}-\theta_{1}\|_{H}^{2/r}\left\|\Lambda^{\alpha}(\theta_{2}-\theta_{1})\right\|_{H}^{2(1-1/r)}\,\mathrm{d}s$
$\leqslant\frac{1}{2}\int_{0}^{t}\left\|\theta_{2}-\theta_{1}\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+C\int_{0}^{t}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}-\theta_{1}\right\|_{H}^{2}\,\mathrm{d}s,$
In (A.14), we apply Hölder’s inequality in the second line, Calderón-Zygmund inequality in the third line, Gagliardo-Nirenberg interpolation inequality in the fourth line and Young’s equality in the last line. Substituting (A.14) into (A.13), we conclude that
$\|\theta_{1}(t)-\theta_{2}(t)\|_{H}^{2}+\int_{0}^{t}\left\|\theta_{1}(s)-\theta_{2}(s)\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant C\int_{0}^{t}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}-\theta_{1}\right\|_{H}^{2}\,\mathrm{d}s.$
It follows from Grönwall’s inequality that $\theta_{1}-\theta_{2}=0.$
In the next example, we will take $\mathcal{C}^{2}:=\mathbb{X}_{\alpha,\alpha/2}\cap L^{4}\left([0,T];H^{1-\alpha,4}(\mathbb{T}^{2})\right)$ and show that $\mathcal{C}^{2}$ is an $\mathcal{I}$-closure of $\mathcal{C}_{0}^{2}.$ Here $H^{1-\alpha,4}(\mathbb{T}^{2})$ is the homogeneous Sobolev space defined by
@@
where $p_{1},p_{2}\in(1,\infty)$ satisfy
$\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}.$
@@
Let $\theta\in\mathcal{C}^{2}$ with $\mathcal{I}(\theta)<\infty$. Then there exists a sequence $\{\theta^{(n)}\}_{n\geqslant 1}\subset\mathcal{C}_{0}^{2}$ such that $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\alpha/2}$ and $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta)$ as $n\rightarrow\infty$.
@@
Since both Riesz transform and $\Lambda^{1-\alpha}$ are defined by Fourier multipliers, it follows from Calderón-Zygmund inequality that, for all $t\in[0,T]$,
$\|u_{\theta}(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}$ $=\|\Lambda^{1-\alpha}R^{\perp}\theta(t)\|_{L^{4}(\mathbb{T}^{2};\mathbb{R}^{2})}=\|R^{\perp}\Lambda^{1-\alpha}\theta(t)\|_{L^{4}(\mathbb{T}^{2};\mathbb{R}^{2})}$
$\leqslant C\|\Lambda^{1-\alpha}\theta(t)\|_{L^{4}(\mathbb{T}^{2})}=C\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}.$
Using Lemma A.5 with $p_{1}=p_{2}=4$, Cauchy-Schwartz inequality, and Calderón-Zygmund inequality, we have
$\quad\|u_{\theta}\theta\|_{L^{2}\left([0,T];H^{1-\alpha}(\mathbb{T}^{2};\mathbb{R}^{2})\right)}^{2}$
$\lesssim\int_{0}^{T}\|u_{\theta}(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}^{2}\|\theta(t)\|_{L^{4}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t+\int_{0}^{T}\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}^{2}\|u_{\theta}(t)\|_{L^{4}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t$
$\lesssim\bigg{(}\int_{0}^{T}\|u_{\theta}(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}^{4}\,\mathrm{d}t\bigg{)}^{1/2}\bigg{(}\int_{0}^{T}\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}^{4}\,\mathrm{d}t\bigg{)}^{1/2}$
$\quad+\bigg{(}\int_{0}^{T}\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}^{4}\,\mathrm{d}t\bigg{)}^{1/2}\bigg{(}\int_{0}^{T}\|u_{\theta}(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}^{4}\,\mathrm{d}t\bigg{)}^{1/2}$
$\lesssim\int_{0}^{T}\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}^{4}\,\mathrm{d}t.$
@@
Applying Lemma A.5 again, we see that $u_{\theta^{(n)}}\theta^{(n)}-u_{\theta}\theta=(u_{\theta^{(n)}}-u_{\theta})\theta^{(n)}+u_{\theta}(\theta^{(n)}-\theta)\rightarrow 0$ in $L^{2}\left([0,T];H^{1-\alpha}(\mathbb{T}^{2};\mathbb{R}^{2})\right)$. Thus
$R^{2,n}:=\Lambda^{-\alpha}\nabla\cdot\left(u_{\theta^{(n)}}\theta^{(n)}-u_{\theta}\theta\right)\rightarrow 0\quad\text{in}\quad L^{2}\left([0,T];H\right).$
@@
The embedding of $H^{1,\frac{1}{1/4+\alpha/2}}(\mathbb{T}^{2})$ into $H^{1-\alpha,4}(\mathbb{T}^{2})$ and the fact that $\frac{1}{4}+\frac{\alpha}{2}+\frac{\alpha}{4}\geqslant\alpha$ imply that the space $\mathcal{C}^{2}$ is larger than $\mathcal{C}_{0}^{2}$ when taking $r=4$ in the definition of $\mathcal{C}_{0}^{2}$.
Appendix B B Large Deviations for Initial Data
In this section, we will verify the large deviations for the initial data of (1.3) under Assumption A. Let $\mu_{\varepsilon,\delta}^{0}:=\mathcal{G}\left(0,\varepsilon Q_{\delta}/2\right)$ be the Gaussian measure on $H^{\alpha-2\beta}(\mathbb{T}^{2})$. Recall that $\mathcal{I}_{0}:H^{\alpha-2\beta}(\mathbb{T}^{2})\rightarrow[0,\infty)$ is defined by (4.2) and $\Lambda_{0}:C_{b}(H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))\times H^{\alpha-2\beta}(\mathbb{T}^{2})\rightarrow\mathbb{R}$ is defined by (4.5) and (4.6).
Proposition B.1.
Assume that the scaling regime (2.7) holds for $\varepsilon,\delta>0,$ then $\{\mu_{\varepsilon,\delta}^{0}\}_{\varepsilon,\delta>0}$ satisfy the large deviations on $H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})$ with speed $\varepsilon^{-1}$ and rate function $\mathcal{I}_{0}$.
@@ Appendix B B Large Deviations for Initial Data
Upper bound: Let $X_{1}^{\varepsilon,\delta}$ be an $H^{\alpha-2\beta}(\mathbb{T}^{2})$-valued random variable with the law $\mu_{\varepsilon,\delta}^{0}$. Then for any $\eta\in(0,1/2)$, it follows from (2.7) that
@@
$\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{\varepsilon}\left\|X_{1}^{\varepsilon,\delta}\right\|^{2}_{H^{\alpha-2\beta}(\mathbb{T}^{2})}\Big{\}}\Big{)}$ $=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{\varepsilon}|k|^{2\alpha-4\beta}\langle e_{k},X_{1}^{\varepsilon,\delta}\rangle^{2}\Big{\}}\Big{)}$ (B.1)
$=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\varepsilon\log(1-\lambda_{\delta,k}^{2}\eta)^{-1/2}$
$\leqslant\varepsilon\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\eta\lambda_{\delta,k}^{2}\leqslant C.$
Thanks to (B.1), it suffices to prove the weak large deviations for $\{\mu_{\varepsilon,\delta}^{0}\}$. Let $F_{0}$ be an arbitrary weakly compact set in $H^{\alpha-2\beta}(\mathbb{T}^{2})$. Using the definitions of $\Lambda_{0}$ and $\mu_{\varepsilon,\delta}^{0}$, a straightforward calculation shows that, for any $\phi\in F_{0}$,
$\int_{H^{\alpha-2\beta}(\mathbb{T}^{2})}\exp\left\{\varepsilon^{-1}\Lambda_{0}\left(\psi,\phi\right)\right\}\mu_{\varepsilon,\delta}^{0}(\mathrm{d}\phi)\leqslant 1.$
It follows that
$\mu_{\varepsilon,\delta}^{0}(F_{0})$ $=\int_{H^{\alpha-2\beta}(\mathbb{T}^{2})}I_{\{\phi\in F_{0}\}}\mu_{\varepsilon,\delta}^{0}(\mathrm{d}\phi)\leqslant\sup_{\phi\in F_{0}}\exp\left\{-\varepsilon^{-1}\Lambda_{0}\left(\psi,\phi\right)\right\}=\exp\left\{-\varepsilon^{-1}\inf_{\phi\in F_{0}}\Lambda_{0}\left(\psi,\phi\right)\right\}.$
Consequently, we have
$\varepsilon\log\mu_{\varepsilon,\delta}^{0}(F_{0})\leqslant-\inf_{\phi\in F_{0}}\Lambda_{0}\left(\psi,\phi\right).$
Since $F_{0}$ is weakly compact in $H^{\alpha-2\beta}(\mathbb{T}^{2})$, we can apply [KL99, Appendix 2, Lemma 3.2] to swap the infimum and supremum to get
$\varepsilon\log\mu_{\varepsilon,m}^{0}(F_{0})\leqslant-\sup_{\psi\in H^{\infty}(\mathbb{T}^{2})}\inf_{\phi\in F_{0}}\Lambda_{0}\left(\psi,\phi\right)=-\inf_{\phi\in F_{0}}\sup_{\psi\in H^{\infty}(\mathbb{T}^{2})}\Lambda_{0}\left(\psi,\phi\right)=-\inf_{\phi\in F_{0}}\left\|\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}.$
Lower Bound: We will take $E=H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})$ and $I=\mathcal{I}_{0}$ in Lemma 2.5 to obtain the lower bound of large deviations. For any $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$, set $\pi_{\varepsilon,\delta}^{0,\phi}:=\mathcal{G}\left(\sqrt{Q_{\delta}}\phi,\varepsilon Q_{\delta}/2\right)$ and $\rho^{0}_{\varepsilon,\delta}:=\frac{\mathrm{d}\pi_{\varepsilon,\delta}^{0,\phi}}{\mathrm{d}\mu_{\varepsilon,\delta}^{0}}$. With the help of Cameron-Martin theorem [DPZ14, Theorem 2.23], the relative entropy can be computed by
$\operatorname{Ent}\left(\pi_{\varepsilon,\delta}^{0,\phi}|\mu_{\varepsilon,\delta}^{0}\right)=\int_{H^{\alpha-2\beta}(\mathbb{T}^{2})}\log\rho^{0}_{\varepsilon,\delta}\,\mathrm{d}\pi_{\varepsilon,\delta}^{0,\phi}=\frac{1}{\varepsilon}\left\|Q_{\delta}^{-1/2}\sqrt{Q_{\delta}}\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}=\frac{1}{\varepsilon}\left\|\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}.$
This leads to (2.5). To verify the weak convergence condition (i) in Lemma 2.5, let $X_{2}^{\varepsilon,\delta}$ be an $H^{\alpha-2\beta}(\mathbb{T}^{2})$-valued random variable with the law $\pi_{\varepsilon,\delta}^{0,\phi}$. Then (2.7) implies
$\mathbb{E}\left\|X_{2}^{\varepsilon,\delta}-\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $=\left\|\sqrt{Q_{\delta}}\phi-\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\mathbb{E}\left\|X_{2}^{\varepsilon,\delta}-\sqrt{Q_{\delta}}\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$
$=\left\|\sqrt{Q}_{\delta}\phi-\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\varepsilon}{2}\operatorname{Tr}\left(Q_{\delta}\right)\rightarrow 0\quad\text{ as }\varepsilon\rightarrow 0.$
Hence $\pi_{\varepsilon,\delta}^{0,\phi}\rightarrow\delta_{\phi}$ weakly in $H_{w}^{\alpha-2\beta}(\mathbb{T}^{2})$. ∎

### v3

@@ Appendix A A Examples of the Weak-strong Uniqueness Classes
This is a direct consequence of [DC06_weak_strong, Theorem 1.1]. ∎
In the next example, we will take $\mathcal{R}^{2}:=\mathbb{X}_{\alpha,\alpha/2}\cap L^{4}\left([0,T];H^{1-\alpha,4}(\mathbb{T}^{2})\right)$ and show that $\mathcal{R}^{2}$ is contained in the $\mathcal{I}$-closure of $\mathcal{C}_{0}^{2}.$ Here $H^{1-\alpha,4}(\mathbb{T}^{2})$ is the homogeneous Sobolev space defined by
@@
where $p_{1},p_{2}\in(1,\infty)$ satisfy $\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}.$
@@
Let $\theta\in\mathcal{R}^{2}$ with $\mathcal{I}(\theta)<\infty$. Then there exists a sequence $\{\theta^{(n)}\}_{n\geqslant 1}\subset\mathcal{C}_{0}^{2}$ such that $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\alpha/2}$ and $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta)$ as $n\rightarrow\infty$.
@@
Since both Riesz transform and $\Lambda^{1-\alpha}$ are defined by Fourier multipliers, it follows from Calderón-Zygmund inequality that, for all $t\in[0,T],\|u_{\theta}(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}\leqslant C\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}.$ Using Lemma A.5 with $p_{1}=p_{2}=4$, Cauchy-Schwarz inequality, and Calderón-Zygmund inequality, we have
$\|u_{\theta}\theta\|_{L^{2}\left([0,T];H^{1-\alpha}(\mathbb{T}^{2};\mathbb{R}^{2})\right)}^{2}\lesssim\int_{0}^{T}\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}^{4}\,\mathrm{d}t.$
@@
Applying Lemma A.5 again, we see that $u_{\theta^{(n)}}\theta^{(n)}-u_{\theta}\theta=(u_{\theta^{(n)}}-u_{\theta})\theta^{(n)}+u_{\theta}(\theta^{(n)}-\theta)\rightarrow 0$ in $L^{2}\left([0,T];H^{1-\alpha}(\mathbb{T}^{2};\mathbb{R}^{2})\right)$. Thus $R^{2,n}:=\Lambda^{-\alpha}\nabla\cdot\left(u_{\theta^{(n)}}\theta^{(n)}-u_{\theta}\theta\right)\rightarrow 0$ in $L^{2}\left([0,T];H\right).$ Then it follows from the proof of Example A.2 that $\mathcal{I}_{dyna}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}_{dyna}(\theta)$. ∎
@@
The embedding of $H^{1,\frac{1}{1/4+\alpha/2}}(\mathbb{T}^{2})$ into $H^{1-\alpha,4}(\mathbb{T}^{2})$ and the fact that $\frac{1}{4}+\frac{\alpha}{2}+\frac{\alpha}{4}\geqslant\alpha$ imply that the space $\mathcal{R}^{2}$ is larger than $\mathcal{C}_{0}^{2}$ when taking $r=4$ in the definition of $\mathcal{C}_{0}^{2}$.
To illustrate the next example in the critical case $\alpha=1/2$, we recall that the BMO space on the $d$-dimensional torus is defined as
[…]
where $m_{B}f=\int_{B}f(x)\,\mathrm{d}x$. $BMO(\mathbb{T}^{d})$ endowed with the norm
$\|f\|_{BMO(\mathbb{T}^{d})}=\sup_{\{B\subset\mathbb{T}^{d}:B\text{ is a ball}\}}\frac{1}{|B|}\int_{B}\left|f-m_{B}f\right|\,\mathrm{d}x+\left|\int_{\mathbb{T}^{d}}f\,\mathrm{d}x\right|$
is a Banach space. We recall that a function $a$ on $\mathbb{T}^{d}$ is called an atom if $a$ is supported on a ball $B$, $\int_{B}a(x)\,\mathrm{d}x=0$ and $\|a\|_{L^{\infty}(\mathbb{T}^{d})}\leqslant\frac{1}{|B|}.$ The Hardy space $\mathcal{H}^{1}\left(\mathbb{T}^{d}\right)$ is defined as
[…]
The norm of $\mathcal{H}^{1}\left(\mathbb{T}^{d}\right)$ is given by $\|f\|_{\mathcal{H}^{1}(\mathbb{T}^{d})}=\inf\left\{\sum_{j\in\mathbb{N}}\left|\lambda_{j}\right|:f=\sum_{j\in\mathbb{N}}\lambda_{j}a_{j}\right\}$, where the infinimum is taken over all the atomic decompositions of $f$.
Lemma A.8 ([CW77]).
$\mathcal{H}^{1}(\mathbb{T}^{d})$ is a Banach space and the dual space of $\mathcal{H}^{1}(\mathbb{T}^{d})$ is $BMO(\mathbb{T}^{d})$.
Lemma A.9.
For any $f\in L^{2}{(\mathbb{T}^{2})},$ there exists a constant $C>0$ such that
$\left\|fR_{i}f\right\|_{\mathcal{H}^{1}(\mathbb{T}^{2})}\leqslant C\|f\|_{L^{2}(\mathbb{T}^{2})}^{2},\quad i=1,2,$
where $R_{i}f$ is the i-th Riesz transform of $f.$
For more details of Lemma A.9, we refer to [CC04, Section 3] and the references therein.
Example A.10.
Suppose that $\alpha=1/2$ and $\beta=\alpha/2+1/4=1/2$. There exists a positive number $D_{\infty}$ such that $\mathcal{C}_{0}^{3}:=\{\theta\in\mathbb{X}_{1/2,1/2}:\theta\in L^{2}\left([0,T];L^{p}(\mathbb{T}^{2})\right),\|R^{\perp}\theta\|_{L^{\infty}([0,T];BMO(\mathbb{T}^{2}))}\leqslant D_{\infty},p>2\}$ is a weak-strong uniqueness class. More precisely, for every $g\in L^{2}([0,T];H)$, assume that $\theta_{1}\in\mathcal{C}_{0}^{3}$ and $\theta_{2}\in\mathbb{X}_{1/2,1/2}$ solve the skeleton equation (1.1) in the sense of Definition 4.1 with the same initial data $\theta_{1}(0)=\theta_{2}(0)\in H^{-1/2}(\mathbb{T}^{2})$, and suppose further that $\theta_{2}$ satisfy the $H^{-1/2}$-energy inequality: for every $t\in[0,T]$,
$\frac{1}{2}\|\theta_{2}(t)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\theta_{2}(s)\|_{H}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{2}(0)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\theta_{2},g\rangle\,\mathrm{d}s.$ (A.3)
Then $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{1/2,1/2}$.
@@
The same conclusion holds in $\mathbb{R}^{2}$ case, which was proved in [Mar08_Weak_strong_critical, Theorem 1.3]. For every $t\in[0,T]$, we expand $\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}$ as
$\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}=-2\langle\Lambda^{-1/2}\theta_{1}(t),\Lambda^{-1/2}\theta_{2}(t)\rangle+\|\theta_{1}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}+\|\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}$ (A.4)
$\quad\int_{0}^{t}\left(2\langle\theta_{1},\theta_{2}\rangle+\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}\rangle\right)\,\mathrm{d}s$ (A.5)
$=-\langle\Lambda^{-1/2}\theta_{2}(t),\Lambda^{-1/2}\theta_{1}(t)\rangle+\langle\Lambda^{-1/2}\theta_{2}(0),\Lambda^{-1/2}\theta_{1}(0)\rangle+\int_{0}^{t}\langle\theta_{2}+\theta_{1},g\rangle\,\mathrm{d}s.$
This can be proved by taking $\Lambda^{-1}\theta^{(n)}_{i}(i=1,2)$ as test functions and passing to the limits. We focus on the convergence of the nonlinear terms. Since $\theta_{1}\in\mathcal{C}_{0}^{3}$, $\nabla\Lambda^{-1}\theta_{1}=R\theta_{1}\in L^{\infty}([0,T];BMO(\mathbb{T}^{2}))$. Moreover, Lemma A.9 implies that $\theta_{2}u_{\theta_{2}}\in L^{1}([0,T];\mathcal{H}^{1}(\mathbb{T}^{2}))$. By the the equivalent characterization of the $\mathcal{H}^{1}$-norm ([Duo01, Theorem 6.4]) and the property of the convolution kernel $\eta_{1/n}$,
$\left|\int_{0}^{t}\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}(\theta_{1}-\theta^{(n)}_{1})\rangle\,\mathrm{d}s\right|=\left|\int_{0}^{t}\langle\theta_{2}u_{\theta_{2}}-(\theta_{2}u_{\theta_{2}})^{(n)},\nabla\Lambda^{-1}\theta_{1}\rangle\,\mathrm{d}s\right|\rightarrow 0.$
The continuous embedding $BMO(\mathbb{T}^{2})\subset L^{\frac{2p}{p-2}}(\mathbb{T}^{2})$ ([Duo01, Corollary 6.12]) implies that $\theta_{1}u_{\theta_{1}}\in L^{2}([0,T];L^{2}(\mathbb{T}^{2})).$ Since $\theta^{(n)}_{2}\rightarrow\theta_{2}$ in $L^{2}([0,T];L^{2}(\mathbb{T}^{2})),$ it follows that
$\bigg|\int_{0}^{t}\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}(\theta_{2}-\theta^{(n)}_{2})\rangle\,\mathrm{d}s\bigg|\rightarrow 0.$
This completes the proof of (A.5). Since (A.5) implies that $\theta_{1}$ is sufficiently regular to be taken as a test function, for any $t\in[0,T],$
@@
$\frac{1}{2}\|\theta_{1}(t)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\theta_{1}(s)\right\|_{H}^{2}\,\mathrm{d}s=\frac{1}{2}\|\theta_{1}(0)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\theta_{1},g\rangle\,\mathrm{d}s.$ (A.6)
According to (A.4), (A.5), (A.6), and (A.3), using Lemma A.8 and Lemma A.9, we obtain
$\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}+2\int_{0}^{t}\left\|\theta_{1}-\theta_{2}\right\|_{H}^{2}\,\mathrm{d}s$
$\leqslant$ $2\int_{0}^{t}\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}\rangle\,\mathrm{d}s$
$=$ $-2\int_{0}^{t}\langle(\theta_{2}-\theta_{1})u_{\theta_{1}},\nabla\Lambda^{-1}(\theta_{2}-\theta_{1})\rangle\,\mathrm{d}s$
$\leqslant$ $2\|(\theta_{2}-\theta_{1})R(\theta_{2}-\theta_{1})\|_{L^{1}([0,T];\mathcal{H}^{1}(\mathbb{T}^{2}))}\|u_{\theta_{1}}\|_{L^{\infty}([0,T];BMO(\mathbb{T}^{2}))}$
$\leqslant$ $C\|\theta_{2}-\theta_{1}\|^{2}_{L^{2}([0,T];H)}\|R^{\perp}\theta_{1}\|_{L^{\infty}([0,T];BMO(\mathbb{T}^{2}))}$
$\leqslant$ $CD_{\infty}\|\theta_{2}-\theta_{1}\|^{2}_{L^{2}([0,T];H)}.$
Hence we can choose a sufficiently small $D_{\infty}$ to deduce that $\theta_{1}-\theta_{2}=0.$